# Identifying Network Structure of Nonlinear Dynamical Systems: Contraction and Kuramoto Oscillators

### v1

@@ II-B Application II: Linear Systems
To provide an intuitive interpretation of the results previously derived we now consider the linear system setting. We remark that a complete and thorough analysis, via observability, of the linear setting can be found in [7]. Specifically, consider a linear system given as
$\Sigma_{\text{lin}}:\begin{cases}\dot{x}=Ax&x(0)=x_{0}\\ y=Cx\end{cases}.$ (28)
Suppose we perturb the system matrix $A$ to evolve under $A+\Delta$ with $C\Delta=0$. Then to satisfy the the LMI (5) with a constant matrix $M(x)=M$, we require that
@@
From Theorem 1 this amounts to requiring that $\alpha(CAC^{\dagger})\leq-\mu$ and $\mathcal{N}(C)\subseteq\mathcal{N}(CA)$. To illustrate the implications, we consider a simple example
@@
This result is fully consistent with the fact that $x_{1}$ is full decoupled from $x_{2}$ and hence the dynamics of $x_{2}$ can be modified freely, as we would expect from our intuition.
@@ III Kuramoto Oscillator Networks
We now focus our attention on the Kuramoto model of coupled oscillators, which is often used in modelling neural systems [15]. In particular, we now study the impact of this model’s contracting dynamics on its identifiability. We apply our framework to prove that a variety of network topologies are indistinguishable from one other. Specifically as in [9, 16] we write the Kuramoto oscillator network dynamics as
$\Sigma_{\text{Kuramoto}}:\begin{cases}\dot{x}=\omega-B\mathcal{A}\sin(B^{\top}x)\\ y=Cx\end{cases}.$ (32)
Let $e$ be the number of edges in the network, then $B\in\mathbb{R}^{n\times e}$ is the incidence matrix encoding the network topology with arbitrary assignment of sink and source nodes, and $\mathcal{A}\in\mathbb{R}^{e\times e}$ is a diagonal matrix of the edge weights (not to be confused with the adjacency matrix). We note that we always define $B$ to encode the incidence matrix for a complete graph with $n$ nodes, and hence, if an edge is not present we set the corresponding edge weight encoded in $\mathcal{A}$ to be zero. Here $\omega$ encodes the natural frequencies of the oscillators, we assume that the oscillator frequencies are non-identical.
@@
We note that in the case of fully connected networks of identical oscillators, all oscillators would synchronize [8] to the common natural frequency $\omega$ and further all connected networks would be indistinguishable as all observations would differ by at most a constant phase shift.
We now consider how to analyze how variations to the topologies alter the measured dynamics of the oscillators. To do so, we define a vector $\delta\in\mathbb{R}^{e}$ of perturbations to the edge weights, and write the corresponding corresponding auxiliary system as
$\begin{cases}\dot{\mathbf{x}}=\omega-B(s\text{diag}(\delta)+\mathcal{A})\sin(B^{\top}\mathbf{x})\\ \mathbf{y}=C\mathbf{x}\end{cases}.$ (33)
@@
$f(\mathbf{x})=-B\mathcal{A}\sin(B^{\top}\mathbf{x})$ (34a)
@@
$\frac{\partial f}{\partial\mathbf{x}}=-B\mathcal{A}\text{diag}(\cos(B^{\top}\mathbf{x}))B^{\top}$ (35a)
@@
We note that (35) are essentially weighted Laplacian matrices and hence $\mathbbm{1}\in\mathcal{N}(\frac{\partial f}{\partial\mathbf{x}})$ and $\mathbbm{1}\in\mathcal{N}(\frac{\partial\Delta}{\partial\mathbf{x}})$, thus in the case of full observations, $C=I$, we can at most expect semicontraction (contraction in the space orthogonal to $\text{span}(\mathbbm{1})$). We also note that the second smallest eigenvalue of the auxiliary systems Jacobian will always be greater than 0 if and only if
@@
encodes the weighted Laplacian of a connected graph [9]. Furthermore, we see that (15) amounts to determining whether
@@
The conditions for two different oscillator networks to contract to each other are analogous to requirements for synchrony found in [9, 17, 8]. This reiterates previous findings in [5] that establish a link between synchrony and the inability to identify network topologies.
To illustrate the implications of making partial measurements of the states of the networks, we now apply our framework to find other indistinguishable topologies when starting with a candidate topology and corresponding measurement scheme. To do so, we consider an example comprised of four oscillators which is displayed in Fig. 2 (Net 1). In mathematical terms, we define the incidence matrix for the complete graph as
$B\coloneq\begin{bmatrix}1&1&1&0&0&0\\ -1&0&0&1&1&0\\ 0&-1&0&-1&0&1\\ 0&0&-1&0&-1&-1\end{bmatrix}$ (38)
@@
Notably, this network is connected with each oscillator connected to its nearest neighbour. Moreover, we consider non-identical natural frequencies given as
@@
$C\coloneq\begin{bmatrix}1&1&0&0\\ 0&0&1&1\end{bmatrix}$. (42)
The measurement scheme defined in (42) indicates that we are measuring neighbouring nodes in the network in an averaging fashion. We now consider new candidate networks that have an edge weight matrix
@@
which confirms (ii). We now consider satisfying (i), evaluating (15), $C\big{(}\frac{\partial f}{\partial\mathbf{x}}+s\frac{\partial\Delta}{\partial\mathbf{x}}\big{)}C^{\dagger}$, we arrive at,
$\frac{1}{2}(a_{13}\cos(\mathbf{x}_{1}-\mathbf{x}_{3})+a_{24}\cos(\mathbf{x}_{2}-\mathbf{x}_{4}))\begin{bmatrix}-1&1\\ 1&-1\end{bmatrix}$ (45)
We remark here that (45) has one negative eigenvalue and the nullspace is simply equal to $\text{span}(\mathbbm{1})$ provided that
@@
Hence we may conclude, that the observations will be contracting in all space except $\text{span}(\mathbbm{1})$. Also, if (46) is satisfied, then if we can find a matrix $M$ to satisfy (47), this matrix will also satisfy (5), as required.
@@
Specifically, we note that with $M=I-\frac{1}{n}\mathbbm{1}\mathbbm{1}^{\top}$ as in [17], we have that (47) (and therefore (5)) is satisfied. We also note that to verify $\mathcal{N}(C)\subseteq\mathcal{N}(C(\frac{\partial f}{\partial\mathbf{x}}+s\frac{\partial\Delta}{\partial\mathbf{x}}))$, it is sufficient to satsify
@@
We remark here that if this condition, (48), is satisfied following Theorem 1 we can use the suggested matrix $M$ with the caveat that we cannot achieve contractivity in the space defined by $\text{span}(\mathbbm{1})$, and thus we have satisfied requirement (i). Furthermore, we note that since there were no constraints enforced on the remaining perturbation weights, we may arbitrarily alter $\delta_{1}$ and $\delta_{6}$, this implies that all networks shown in Fig. 2 display observations that differ by only a constant phase shift.
The conditions outlined in (46) and (48) are challenging to verify due to the dependence on the state, $\mathbf{x}$, and the parametrization variable, $s$. However, we now outline alternative sufficient conditions for satisfying (46) and (48). To do so, we introduce the following definition similar to those in [8, 17].
@@
Notably, we see that if the oscillator states are $\pi/2$-phase cohesive, i.e. $\mathbf{x}\in\Theta(\frac{\pi}{2})~\forall t\geq 0$, then (46) is satisfied for any positive edge weights $a_{13}$ and $a_{24}$. We note that remaining phase cohesive has been extensively studied as it has a close relationship to achieving synchrony, we point the reader to [8] for a thorough discussion. Typically, the conditions to remain phase cohesive are conservative estimates on the maximum difference between the oscillator frequencies, $\omega$, for instance one such condition from [8] is for
@@
We additionally note here that this condition would not be directly applicable to the disconnected topology in Fig. 2 (Net 4) as $\lambda_{2}(B(\text{diag}(\delta)+\mathcal{A})B^{\top})=0$. However, we do numerically observe that this topology is indeed indistinguishable from the others (see Section III-A).
Now in order to satisfy (48) it suffices to ensure that there is some symmetry, in the sense that permuting the pair of labels 1 and 3 with 2 and 4 will lead to identical difference dynamics, $\mathbf{x}_{1}-\mathbf{x}_{3}$ and $\mathbf{x}_{2}-\mathbf{x}_{4}$, for each of the oscillator pairs. We note that this is trivially achievable if all the oscillator frequencies and edge weights are identical; if not, certain selections of similar frequencies and edge weights are sufficient.
III-A Simulations
In order to validate our findings we simulate all four networks from Fig. 2 and plot the corresponding measurements $\mathbf{y}_{1}(t)$ and $\mathbf{y}_{2}(t)$ in Fig. 3.
Figure 3: Comparison of different Kuramoto networks displayed in Fig. 2. When beginning with the same initial condition (Left column), all four network topologies’ measurements are identical to each other over the entire measurement period and have no phase shifts between them. When the networks start with different initial conditions (Right column) the states differ by a constant phase shift.
We see, as expected, in Fig. 3 that when the networks are started with all the same initial conditions (Fig. 3 (Left)) they do not display any phase shift between them, and are identical for all time, implying that the networks are all indistinguishable from one another. Moreover, when the initial conditions are different (Fig. 3 (Right)), the networks have measurements that share a common waveform and only differ by a fixed phase shift. We note that, despite differing by a phase shift the networks are still indistinguishable as measuring the behavior past the transient dynamics period we see that the functional form of the measurements are identical. Remarkably, we also highlight that in the partial measurement setting it is possible to mischaracterize connected networks as disconnected networks (Fig. 2 (Net 4)) as illustrated by this example. Moreover, we note that this example also emphasizes once more that synchrony yields challenges in identification.
Displayed in Fig. 4 the oscillators were simulated with random natural frequencies, $\omega$, and random initial conditions, $x_{0}$, which do not necessarily adhere to the requirement (48), we notice that despite this, the networks still observe similar behavior as in Fig. 3. This emphasizes that the conditions are sufficient but not necessary.
@@ III-A Simulations
In summary, in order to consider whether other candidate network topologies are indistinguishable from a hypothesis network it it suffices to establish satisfaction of requirements (i) and (ii) (i.e., applying Theorem 1 and Corollary 2) outlined at the end of Section II.
IV Conclusions and Future Work
In this paper we have established a series of results that describe how contractivity (in a single dynamical system and also in the observable space) can result in challenges in the identification of network topologies. Namely, we provide sufficient conditions for observed trajectories of different nonlinear systems to contract to one another by leveraging the contraction theory framework and establishing a notion of contractivity in the observable space. Particularizing our results to Kuramoto oscillator networks we established that synchrony, phase cohesiveness, and symmetry play a key role in determining whether different topologies are indistinguishable from one another. We anticipate the framework proposed will find useful applications to various other models of neural dynamics in order to study how different implementations of brain networks or neural circuits exhibit the same function and can help guide new techniques for the inference of the topologies in these network models.

### v2

@@ II-B Application II: Linear Systems
To build some intuition, we now consider linear systems. A complete and thorough analysis, via observability, of the linear setting can be found in [6]. Specifically, consider a linear system given as
$\Sigma_{\text{lin}}:\begin{cases}\dot{x}=Ax,&x(0)=x_{0}\\ y=Cx.\end{cases}$ (28)
Suppose we perturb the system matrix $A$ to evolve under $A+\Delta$ with $C\Delta=0$. Then to satisfy (5) with a constant matrix $M(x)=M$, we need
@@
From Theorem 1 we require $\alpha(CAC^{\dagger})\leq-\mu$ and $\mathcal{N}(C)\subseteq\mathcal{N}(CA)$. Consider the following example:
@@
This result is fully consistent with the fact that $x_{1}$ is fully decoupled from $x_{2}$ and hence the dynamics of $x_{2}$ can be modified freely, as expected.
@@ III Kuramoto Oscillator Networks
We now focus our attention on the Kuramoto model of coupled oscillators, which is often used in modeling neural systems [15]. We study the impact of this model’s contracting dynamics on its identifiability. We apply our framework to prove that a variety of network structures are indistinguishable from each other. As in [8, 17] we express the Kuramoto oscillator network dynamics as
$\Sigma_{\text{Kuramoto}}:\begin{cases}\dot{x}=\omega-B\mathcal{A}\sin(B^{\top}x),\\ y=Cx.\end{cases}$ (32)
Let $e$ be the number of edges in the network, then $B\in\mathbb{R}^{n\times e}$ is the incidence matrix encoding the network structure with arbitrary assignment of sink and source nodes, and $\mathcal{A}\in\mathbb{R}^{e\times e}$ is a diagonal matrix of the edge weights (not to be confused with the adjacency matrix). We always define $B$ to encode the incidence matrix for a complete graph with $n$ nodes, and hence, if an edge is not present we set the corresponding edge weight encoded in $\mathcal{A}$ to be zero. Here $\omega$ encodes the non-identical natural frequencies of the oscillators.
@@
In the case of fully connected networks of oscillators with identical natural frequencies, all oscillators would synchronize [5] to the common natural frequency $\omega$ and all connected networks would be indistinguishable as all observations would differ by at most a constant phase shift.
We now consider how to analyze how variations to the network structure alter the measured dynamics of the oscillators. We define $\delta\in\mathbb{R}^{e}$ to be a perturbation to the edge weights, then the corresponding auxiliary system is
$\begin{cases}\dot{\mathbf{x}}=\omega-B(s\text{diag}(\delta)+\mathcal{A})\sin(B^{\top}\mathbf{x}),\\ \mathbf{y}=C\mathbf{x}.\end{cases}$ (33)
@@
$f(\mathbf{x})=-B\mathcal{A}\sin(B^{\top}\mathbf{x}),$ (34a)
@@
$\frac{\partial f}{\partial\mathbf{x}}=-B\mathcal{A}\text{diag}(\cos(B^{\top}\mathbf{x}))B^{\top},$ (35a)
@@
The Jacobians (35) are weighted Laplacian matrices and hence $\mathds{1}\in\mathcal{N}(\frac{\partial f}{\partial\mathbf{x}})$ and $\mathds{1}\in\mathcal{N}(\frac{\partial\Delta}{\partial\mathbf{x}})$, thus in the case of full observations $C=I$ we can at most expect semicontraction (i.e., contraction in the space orthogonal to $\text{span}(\mathds{1})$). The second smallest eigenvalue of the auxiliary systems’ Jacobian will always be greater than 0 if and only if
@@
encodes the weighted Laplacian of a connected graph [8]. Then, (15) amounts to determining whether
@@
The conditions for two different oscillator networks to contract to each other are analogous to requirements for synchrony found in [8, 4, 5]. This reiterates previous findings in [3] that establish a link between synchrony and the inability to identify network structures.
We now apply our framework to find other indistinguishable networks when starting with a candidate network and corresponding measurement scheme. We consider an example comprised of four oscillators which is displayed in Fig. 2 (Net 1). We define the incidence matrix for the complete graph as
$B=\begin{bmatrix}1&1&1&0&0&0\\ -1&0&0&1&1&0\\ 0&-1&0&-1&0&1\\ 0&0&-1&0&-1&-1\end{bmatrix}$ (38)
@@
This network is connected with each oscillator connected to its nearest neighbor. We consider non-identical natural frequencies given as
@@
$C=\begin{bmatrix}1&1&0&0\\ 0&0&1&1\end{bmatrix}$. (42)
The measurement scheme defined in (42) indicates that we are measuring neighboring nodes in an averaging fashion. Now, consider new networks with an edge weight matrix
@@
which confirms (ii). We now consider satisfying (i), evaluating (15), i.e., $C\big(\frac{\partial f}{\partial\mathbf{x}}+s\frac{\partial\Delta}{\partial\mathbf{x}}\big)C^{\dagger}$ we arrive at
$\frac{1}{2}(a_{13}\cos(\mathbf{x}_{1}-\mathbf{x}_{3})+a_{24}\cos(\mathbf{x}_{2}-\mathbf{x}_{4}))\begin{bmatrix}-1&1\\ 1&-1\end{bmatrix}$. (45)
Note that (45) has one negative eigenvalue and the nullspace is simply equal to $\text{span}(\mathds{1})$ provided that
@@
Hence we may conclude, that the observations will be contracting in all space except $\text{span}(\mathds{1})$. Also, if (46) is satisfied, then if we can find a matrix $M$ to satisfy (47), this matrix will also satisfy (5), as required.
@@
Specifically, with $M=I-\frac{1}{n}\mathds{1}\mathds{1}^{\top}$ as in [4], (47) (and therefore (5)) is satisfied. To verify $\mathcal{N}(C)\subseteq\mathcal{N}(C(\frac{\partial f}{\partial\mathbf{x}}+s\frac{\partial\Delta}{\partial\mathbf{x}}))$, it is sufficient to satisfy
@@
If (48) is satisfied following Theorem 1 we can use the suggested matrix $M$ with the caveat that we cannot achieve contractivity in the space defined by $\text{span}(\mathds{1})$, and thus we have satisfied requirement (i). Furthermore, since there were no constraints enforced on the remaining perturbation weights, we may arbitrarily alter $\delta_{1}$ and $\delta_{6}$, this implies that all networks shown in Fig. 2 display observations that differ by only a constant phase shift.
The conditions outlined in (46) and (48) are challenging to verify due to the dependence on the state $\mathbf{x}$ and the parametrization variable $s$. However, we now outline alternative sufficient conditions for satisfying (46) and (48). We introduce the following definition similar to those in [5, 4].
@@
Notably, if the oscillator states are $\pi/2$-phase cohesive, i.e., $\mathbf{x}\in\Theta(\frac{\pi}{2})~\forall t\geq 0$, then (46) is satisfied for any positive edge weights $a_{13}$ and $a_{24}$. Remaining phase cohesive has been extensively studied as it has a close relationship to achieving synchrony, we point the reader to [5] for a thorough discussion. Typically, the conditions to remain phase cohesive are conservative estimates on the maximum difference between the oscillator frequencies $\omega$, for instance, one such condition from [5] is for
@@
This condition would not be directly applicable to the disconnected topology in Fig. 2 (Net 4) as $\lambda_{2}(B(\text{diag}(\delta)+\mathcal{A})B^{\top})=0$. However, in Section IV we observe that this network is still indistinguishable from the others.
To satisfy (48) it suffices to ensure that there is some symmetry, in the sense that permuting the pair of labels 1 and 3 with 2 and 4 will lead to identical difference dynamics $\mathbf{x}_{1}-\mathbf{x}_{3}$ and $\mathbf{x}_{2}-\mathbf{x}_{4}$ for each of the oscillator pairs. This is trivially achievable if all the oscillator frequencies and edge weights are identical; if not, certain selections of similar frequencies and edge weights are sufficient.
IV Simulations
To validate our findings we simulate all four networks from Fig. 2 and plot the corresponding measurements $y_{1}(t)$ and $y_{2}(t)$ in Fig. 3. For Fig. 2 (Right), we select initial conditions for nodes 1 and 2 randomly. In order to satisfy (48), node 4 has the same initial condition as node 1; node 3 has the same initial condition as node 2.
Figure 3: Comparison of different Kuramoto networks displayed in Fig. 2. When beginning with the same initial condition (Left column), all four network structures’ measurements are identical to each other over the entire measurement period and have no phase shifts between them. When the networks start with different initial conditions (Right column) the states differ by a constant phase shift.
We see, as expected, in Fig. 3 that when the networks are started with all the same initial conditions (Fig. 3 (Left)) they do not display any phase shift between them, and are identical for all time, implying that the networks are all indistinguishable from one another. When the initial conditions are different (Fig. 3 (Right)), the networks have measurements that share a common waveform and only differ by a fixed phase shift. Despite differing by a phase shift the networks are still indistinguishable as measuring the behavior past the period of transient dynamics the functional form of the measurements are identical. Remarkably, in the partial measurement setting it is possible to mischaracterize connected networks as disconnected networks (Fig. 2 (Net 4)) as illustrated by this example.
Displayed in Fig. 4 the oscillators were simulated with random natural frequencies $\omega$ and random initial conditions $x_{0}$ which do not necessarily adhere to the requirement (48). Despite this, the networks still observe similar behavior as in Fig. 3. This emphasizes that the conditions established are sufficient but not necessary.
@@ IV Simulations
In summary, in order to consider whether other candidate network structure are indistinguishable from a hypothesis network it suffices to establish satisfaction of requirements (i) and (ii) (i.e., applying Theorem 1 and Corollary 2) outlined at the end of Section II.
V Conclusions and Future Work
In this paper we have established a series of results that describe how contractivity (in a single dynamical system and in the observable space) can result in challenges in the identification of network topologies. We provide sufficient conditions for observed trajectories of different nonlinear systems to contract to one another by leveraging the contraction theory framework and establishing a notion of contractivity in the observable space. Particularizing our results to Kuramoto oscillator networks we established that synchrony, phase cohesiveness, and symmetry play a key role in determining whether different topologies are indistinguishable from one another. We anticipate the framework proposed will find useful applications to various other models of neural dynamics in order to study how different implementations of brain networks or neural circuits exhibit the same function and can help guide new techniques for the inference of the structures in these network models.